Tungsten crystallizes according to BCC unit cell where the edge length of the unit cell is 316.52pm. Use this information to address the following: Use the unit cell edge length to calculate the atomic radius W. How does your answer compare to the known literature value (NOTE: The literature value is 139pm).
Answer #1

FOR BCC:

a*sqrt(3) = 4r

Here:

a is edge length of cube

r is radius of atom

use:

a*sqrt(3) = 4r

316.52 * 1.732 = 4*r

r = 137 pm

% difference from literature value = (139-137)*100/137

= 1.46 %

Answer:

radius = 137 pm

% difference from literature value =1.46 %